Laser light of wavelength 480 nm is incident on a circular aperture which has a diameter of 0.011 mm.
A diffraction pattern is observed on a screen which is placed 94 cm from the aperture.

Give your answer to at least three significant figures. Answer must be accurate to 1%.

diffraction angle, θ, of the first diffraction minimum:
3.051662  degrees

You are correct.

1) What is the distance, on the screen, from the center of the central bright spot to the first dark ring?
Answer #1

Given

Wavelength λ = 480 nm = 480 * 10^-9 m

Diameter of the aperture d = 0.011 mm = 0.011* 10^-3 m

Distance between screen and the aperture D = 94 cm = 0.94 m

Known

For the circular aperture diffraction m = 1.220 for first diffraction minimum

Solution

1)

The condition for first diffraction minimum is

dsinθ = mλ

sinθ = mλ/d

sinθ = 1.220 * 480 * 10^-9  /0.011 * 10^-3

θ = sin-1(0.0532)

θ = 3.05o

2)

For the dark ring the distance from center

y = D tanθ

y = 0.94 * tan 3.05

y = 0.0501 cm

I hope help you.